# Verification of Geometric Robustness of Neural Networks via Piecewise Linear Approximation and Lipschitz Optimisation

### v1

@@ 6 Conclusions
Despite the positive results brought by our piecewise linear approximation method, a few other interesting topics are worth further exploration. First, it is challenging to obtain the optimal piecewise linear constraints via (6). To get a good set of piecewise linear constraints, our current method (7) requires to obtain a good heuristic partition of the domain $\mathcal{B}_{1},\ldots,\mathcal{B}_{q}$. It will be interesting to further investigate and quantify the suboptimality of the solution from (7). Second, the number of piecewise linear segment $q$ is a hyperparameter in our framework. A larger value $q$ leads to a better approximation of the pixel value function in theory; however, this also results in more linear constraints for the verification problem in practice. Future work will investigate how to choose a good value of $q$ based on the curvature of of the pixel value function.

### v2

@@ 6 Conclusions
Despite the positive results brought by our piecewise linear approximation method, further topics deserve further exploration. Firstly, it remains challenging to obtain the optimal piecewise linear constraints via (6). To get a good set of piecewise linear constraints, our current method (7) requires to obtain a good heuristic partition of the domain $\mathcal{B}_{1},\ldots,\mathcal{B}_{q}$. It will be interesting to further investigate and quantify the suboptimality of the solution from (7). Second, the number of piecewise linear segment $q$ is a hyperparameter in our framework. A larger value $q$ leads to a better approximation of the pixel value function in theory; however, this also results in more linear constraints for the verification problem in practice. Future work will investigate how to choose a good value of $q$ based on the curvature of of the pixel value function.
Acknowledgements
Ben Batten was funded by the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence. Alessandro De Palma was supported by the “SAIF" project, funded by the “France 2030” government investment plan managed by the French National Research Agency, under the reference ANR-23-PEIA-0006. Alessio Lomuscio was supported by a Royal Academy of Engineering Chair in Emerging Technologies.
Appendix
Appendix A Linear optimisation over sub-domains
Our discussion below focuses on $q=2$ in (6), and with this choice, we have already found promising improvements in our experiments (see the main text). With this constraint we can find suboptimal piecewise bounds by solving two independent linear optimisation problems, where each problem is applied over a subset of the piecewise domain, divided at a given sample point, $n$. We name the parameter sub-spaces divided by $n$, $\boldsymbol{\kappa}_{\mathfrak{1}}$, and $\boldsymbol{\kappa}_{\mathfrak{2}}$ where $\boldsymbol{\kappa}_{\mathfrak{1}},\boldsymbol{\kappa}_{\mathfrak{2}}\subset\mathcal{B}$. Expressing (6) in this way gives
$\min_{\underline{w}_{1},\underline{b}_{1}}\quad$ $\frac{1}{N}\sum_{i=1}^{n}\left(\mathcal{G}_{u,v}(\kappa_{i})-\{\underline{w}_{1}^{{\mathsf{T}}}\kappa_{i}+\underline{b}_{1}\}\right)$ (16a)
subject to $\{\underline{w}_{1}^{{\mathsf{T}}}\kappa_{i}+\underline{b}_{1}\}\leq\mathcal{G}_{u,v}(\kappa_{i}),\quad i=1,\ldots,N,$
$\min_{\underline{w}_{2},\underline{b}_{2}}\quad$ $\frac{1}{N}\sum_{i=n}^{N}\left(\mathcal{G}_{u,v}(\kappa_{i})-\{\underline{w}_{2}^{{\mathsf{T}}}\kappa_{i}+\underline{b}_{2}\}\right)$ (16b)
subject to $\{\underline{w}_{2}^{{\mathsf{T}}}\kappa_{i}+\underline{b}_{2}\}\leq\mathcal{G}_{u,v}(\kappa_{i}),\quad i=1,\ldots,N.$
In (16a) and (16b) we optimise the area over over only the sample points within a piece’s domain, $\boldsymbol{\kappa}_{\mathfrak{1}}$, or $\boldsymbol{\kappa}_{\mathfrak{2}}$; however, we enforce the constraints at every sample point. By doing this we guarantee convexity of our piecewise constraints. We develop a heuristic to determine the sample point, $n$, at which we split $\mathcal{B}$ based on the error between sampled points and optimal linear bounds
$\underline{n}=\max_{i=1,\ldots,N}\left(\{\overline{w}^{{\mathsf{T}}}\kappa_{i}+\overline{b}\}-\mathcal{G}_{u,v}(\kappa_{i})\right),$ (17)
where $\underline{n}$ is the splitting point for the lower bound. We calculate $\overline{n}$ correspondingly using the lower linear bound. There exists a splitting point, $n$, that would produce optimal piecewise bounds, but finding it is infeasible. Our approach permits splitting either the lower or upper linear bound into a piecewise linear bound. To select the bound to be split we perform piecewise bound generation on both candidates, using the splitting points, $\underline{n}$ and $\overline{n}$, respectively for lower and upper bounds. In each case we select the bound combination that results in the smallest enclosed region.
Appendix B Details of branch-and-bound procedure
With our unsound constraints, our method closely follows that of [1], with the important exception that we treat our single piecewise bound as two, separate linear bounds with domains, $\boldsymbol{\kappa}_{\mathfrak{1}}$, and $\boldsymbol{\kappa}_{\mathfrak{2}}$. We first define a function, $\underline{f}_{u,v}$, to track the violation of a bound by the pixel value function, $\mathcal{G}_{u,v}$. In the case that the lower bound is piecewise, we will maximise $\underline{f}_{u,v}$ twice over $\boldsymbol{\kappa}_{\mathfrak{1}}$, and $\boldsymbol{\kappa}_{\mathfrak{2}}$, and $\overline{f}_{u,v}$ once over $\mathcal{B}$. Maximisation of $\underline{f}_{u,v}$ is done via a branch-and-bound Lipschitz procedure. Algorithm 1 shows a simplified version of the implementation we use. For each instance of $\underline{f}_{u,v}(\kappa)\text{ where }\kappa\in\boldsymbol{\kappa}_{\mathfrak{1}}$, we first approximate the Lipshitz constant, $L_{i}$, and use it to bound $\underline{f}_{u,v}$
$\centering\underline{f}_{\text{bound}}=L_{i}\frac{\boldsymbol{\kappa}_{\mathfrak{1}}}{2}+\left(\underline{f}_{u,v}(\kappa_{i})\right),\@add@centering$ (18)
[…]
Algorithm 1 Branch-and-bound Lipschitz Optimisation Procedure
Input: $\underline{f}_{u,v}$, $\mathcal{B}$, $\epsilon$, $n$, $N$
Output: $\xi^{*}$
5:     for $i\leftarrow 1$ to $N$ do
7:           $\mathcal{B}_{i,i+N}=\text{split}(\mathcal{B}_{i})$.
11:        end if
12:     end for
13:  end while
14:  return $\xi^{*}=\max_{i=1,\ldots,N}\underline{f}_{{\max},i}$.